Δ(GPE) = GPE_{f} − GPE_{i}

\frac{Δ(GPE)}{Δheight} = \frac{GPE_{f} − GPE_{i}}{height_{f} − height_{i}}

Δ(SPE) = SPE_{f} − SPE_{i}

\frac{Δ(SPE)}{Δlength} = \frac{SPE_{f} − SPE_{i}}{length_{f} − length_{i}}